### v1

@@ Abstract
A generalized Nash equilibrium problem (GNEP) in Banach space consists of $N>1$ optimal control problems with couplings in both the objective functions and, most importantly, in the feasible sets. We address the existence of equilibria for convex GNEPs in Banach space. We show that the standard assumption of lower semicontinuity of the set-valued constraint maps – foundational in the current literature on GNEPs – can be replaced by graph convexity or the so-called Knaster–Kuratowski–Mazurkiewicz (KKM) property. Lower semicontinuity is often essential for obtaining upper semicontinuity of best response maps, crucial for the existence theory based on Kakutani–Fan fixed-point arguments. However, in function spaces or PDE-constrained settings, verifying lower semicontinuity becomes much more challenging (even in convex cases), whereas graph convexity, for example, is often straightforward to check. Our results unify several existence theorems in the literature and clarify the structural role of constraint maps. We also extend Rosen’s uniqueness condition to Banach spaces using a multiplier bias framework. Additionally, we present a geometric counterpart to our analytic framework using preference maps. This geometric is intended as a complement to, rather than a replacement for, the analytic theory developed in the main body of the paper.
@@ 1 Introduction
Game theory provides a mathematical framework for studying strategic interactions among decision makers. Games can be cooperative, if enforceable agreements are allowed, or noncooperative [harsanyi_general_1988], and since the seminal work of von Neumann and Morgenstern [von_neumann_theory_2007] and Nash’s foundational equilibrium concepts [nash_equilibrium_1950, nash_non-cooperative_1951], equilibrium analysis has become a central paradigm in economics, engineering, and the social sciences. A generalized Nash equilibrium problem (GNEP) extends the classical setting by allowing each player’s feasible set to depend on the strategies of others. This formulation captures shared or coupled constraints, which arise naturally in models such as electricity markets [hobbs_nash-cournot_2007, jing-yuan_spatial_1999], coupled hydrogen and electricity markets [dvurechensky2024cournotnashmodelcoupledhydrogen], environmental management [breton_game-theoretic_2006], and resource allocation [roughgarden_algorithmic_2010].
@@
Most classical results on GNEPs are finite-dimensional. However, important applications such as the coupling energy markets to physical transport of energy carriers (see, e.g., [dvurechensky2024cournotnashmodelcoupledhydrogen] and the references therein) lead to GENEPs with partial differential equations (PDEs) modeling the transport. But infinite-dimensional formulations, particularly those constrained by PDEs, remain less developed due to compactness and regularity issues. While some works analyze PDE-constrained Nash equilibrium problems via optimality systems [roubicek_nash_2007, roubicek_noncooperative_1999, ramos_nash_2002, ramos_pointwise_2002, MR3032892], genuine generalized frameworks for multiple players have been addressed mainly under jointly convex assumptions [hintermuller_generalized_2015, hintermuller_generalized_2024, gahururu_generalized_2022, gahururu_risk-neutral_2023, hintermueller_pde-constrained_2013, kanzow_multiplier-penalty_2019]. These studies typically require lower semicontinuity of the constraint maps to invoke Kakutani–Fan-type fixed-point theorems. While this is a standard assumption in abstract analysis, it is in general not very often verifiable in function space settings.
@@
This paper revisits the existence theory for convex GNEPs in Banach spaces to clarify the trade-off between structural assumptions, such as convexity and regularity of set-valued constraint maps. We show that equilibrium existence can be ensured under graph-convexity or the KnasterKuratowski-Mazurkiewicz (KKM) property without assuming lower semicontinuity. These geometric conditions are often much easier to verify in infinite-dimensional or PDE-constrained contexts and thus broaden the class of problems where equilibrium existence can be rigorously established. This relaxes one of the most restrictive assumptions in the standard theory while retaining the classical framework as a special case.
@@
Beyond existence, we study several structural subclasses of convex games. We show that Rosen’s uniqueness condition for variational equilibria [rosen_existence_1965] extends naturally to Banach spaces, and introduce the notion of multiplier bias as a mechanism for equilibrium selection. We also discuss jointly convex and potential games as key examples illustrating the trade-off between structural assumptions and regularity requirements.
@@
Compared with the existing literature on convex GNEPs and QVIs [facchinei_generalized_2007, harker_generalized_1991, pang_quasi-variational_2005, chan_generalized_1982], the novelty is not in a single new fixed-point theorem, but in the way structural assumptions on the constraint maps and the objective functionals can be traded against regularity assumptions, with an eye on PDE-constrained applications. The paper thus complements, rather than overlaps with, the jointly convex theory in [hintermuller_generalized_2015, kanzow_multiplier-penalty_2019] and with equilibrium results in function spaces [mas-colell_equilibrium_1991].
1.1 Outline
@@
Section 2 introduces the mathematical structure of generalized games in Banach spaces. Section 3 revisits classical existence results and highlights the role of lower semicontinuity in the proofs based on the Kakutani’s fixed-point theorem. Section 4 discusses the trade-off between structure and regularity, focusing on jointly convex and potential games. Section 5 establishes equilibrium existence under graph-convexity and KKM assumptions. Section 6 extends Rosen’s uniqueness result and interprets it as a selection mechanism through multiplier bias. Section 7 then presents the geometric counterpart of our setting and discusses when and how the main analytic ideas extend to preference maps.
@@ 2 Generalized finite games in Banach spaces
The strategy set of each player is a fixed Banach space $X_{i}$, likely of infinite dimension, with its topological dual denoted by $X_{i}^{*}$. Representing their private (or individual) constraints, each player observes a nonempty, closed and convex subset $X_{i}^{\rm{ad}}\subset X_{i}$. We write $X\coloneqq X_{1}\times\cdots\times X_{N}$, analogously for $X^{*}$, and $X^{\rm{ad}}\coloneqq X_{1}^{\rm{ad}}\times\cdots\times X_{N}^{\rm{ad}}$ to denote the full strategy space, its topological dual, and the admissible strategy set, respectively. In game theory, it is common to use the notation $x_{-i}$ to denote the bundle of N1N-1 strategies containing all strategies of a given bundle $x\in X$ except that of player $i.$ Moreover, we write $x=(x_{i},x_{-i})$ when emphasis on player $i$’s strategy is needed, but without changing the original order of the components of $x$. Consequently, we let $X_{-i}\coloneqq X_{1}\times\cdots\times X_{i-1}\times X_{i+1}\times\cdots\times X_{N}$ (and the corresponding $X_{-i}^{\rm{ad}}$) and say that $x_{-i}\in X_{-i}.$ In earlier literature on Nash games the bundle $x_{-i}$ is sometimes referred to as an $i$-incomplete combination [harsanyi_general_1988].
@@
$\begin{cases}\text{Given }x_{-i}\in X_{-i}^{\rm{ad}};\\ \text{Minimize }\mathcal{J}_{i}(x_{i},x_{-i})\ \text{subject to }x_{i}\in X_{i}^{\rm{ad}}\cap{\bf X}_{i}(x_{-i}).\end{cases}$ ($P_{i}$)
@@
The simplest situation in which generalized games appear is when a constraint is shared among players. This is the case when, for example, a regulatory individual or agency determines that the strategy bundle $x$ must belong to a certain closed set $\mathcal{C}\subset X.$ This leads to non-constant constraint maps
@@
where $\pi_{i}$ denotes the projection onto $X_{i}$. In Figure 1 we illustrate three different situations regarding the convexity of $\mathcal{C}$ and ${\bf X}_{i}(\cdot).$
[…]
(a) $\mathcal{C}$ and ${\bf X}_{i}(\cdot)$ convex
[…]
(b) $\mathcal{C}$ non-convex, ${\bf X}_{i}(\cdot)$ convex
[…]
(c) $\mathcal{C}$ and ${\bf X}_{i}(\cdot)$ non-convex
Figure 1: Convexity of $\mathcal{C}$ versus convexity of ${\bf X}_{i}(x_{-i})$ for a two-player game at $x=(x_{1},x_{2}).$
We are interested in the concept of Nash equilibrium [nash_equilibrium_1950, nash_non-cooperative_1951] as the notion of solution for a generalized game. A game is at a Nash equilibrium if no player can (or have an incentive to) deviate from their current strategy.
@@
With a very simple finite dimensional example we illustrate that constant versus non-constant constraint maps may change the set of equilibria drastically.
Example 2.3.
Suppose two companies $I=\{1,2\}$ choose quantities $x_{i}$ to produce and sell a homogeneous product on the same market. The selling price is given by the linear inverse demand function
$P(x_{1},x_{2})=\eta-p(x_{1}+x_{2}),\qquad\eta,p>0,$
and the unit production cost of firm $i$ satisfies $0<c_{i}<\eta$. The loss of firm $i$ is
$J_{i}(x_{1},x_{2})=-P(x_{1},x_{2})x_{i}+c_{i}x_{i}=-x_{i}\big(\eta-c_{i}-p(x_{1}+x_{2})\big).$
Let $X_{i}=\mathbb{R}$ and $X_{i}^{\rm{ad}}=\mathbb{R}_{+}$ for $i=1,2$. If the constraint maps are constant, ${\bf X}_{i}(\cdot)\equiv X_{i}^{\rm{ad}}$, each firm solves
$\min_{x_{i}\geqslant 0}J_{i}(x_{i},x_{-i}),$
and the (unconstrained) best response of player $i$ is
$\Phi_{i}(x_{-i})=\frac{\eta-c_{i}-px_{-i}}{2p}.$
The Nash equilibrium $(x_{1}^{\ast},x_{2}^{\ast})$ is uniquely determined as
$x_{1}^{\ast}=\Phi_{1}(x_{2}^{\ast}),\qquad x_{2}^{\ast}=\Phi_{2}(x_{1}^{\ast}).$
Now introduce a shared capacity constraint
$x_{1}+x_{2}\leqslant C$
for some constant $C>0$. Then the feasible set of player $i$ becomes ${\bf X}_{i}(x_{-i})=[0,\,C-x_{-i}]$. Thus, each best response is the truncation of $\Phi_{i}$ to $0,C-x_{-i}]$. If $C$ is large enough so that the unconstrained equilibrium $(x_{1}^{\ast},x_{2}^{\ast})$ satisfies $x_{1}^{\ast}+x_{2}^{\ast}<C$, the equilibrium is unique and unchanged. However, if $C$ is small, the intersection of the unconstrained best-response curves lies outside the feasible region. In that regime, the equilibrium set becomes a full segment on the boundary $\{(x_{1},x_{2})\in\mathbb{R}_{+}^{2}:x_{1}+x_{2}=C\}$, and the game admits infinitely many (indeed uncountably many) generalized Nash equilibria. For an illustration, see Figure 2. ∎
$x_{1}$$_{2}xx^{\ast}$
(a) Large $C$: unique equilibrium.
$x_{1}$$_{2}xx^{\ast}$GNE
(b) Small $C$: continuum of equilibria.
Figure 2: In the $(x_{1},x_{2})$–plane, player 1’s best response is $x_{1}=\Phi_{1}(x_{2})$ (blue) and player 2’s best response is $x_{2}=\Phi_{2}(x_{1})$ (red). For large production capacity (left), the intersection $x^{\ast}$ is feasible and unique. For a small production capacity (right), $x^{\ast}$ becomes infeasible and a continuum of GNE lie on the boundary of the feasible set
In addition to non-uniqueness, the example above points out that in the general case we cannot expect that a single factor will be enough to guarantee that equilibria exist. In games with nonsmooth objectives, the issue of compacteness becomes essential. The next example can be found in [mas-colell_equilibrium_1991].
Example 2.4.
Consider a two-player game $I=\{1,2\}$ in which each player chooses a continuous nonnegative function on $[0,1]$. For $i=1,2$, let
$X_{i}=\{x\in C([0,1])\,:\,x(t)\geqslant 0\ \text{for all }t\in[0,1]\},\qquad X_{i}^{\mathrm{ad}}=\{x_{i}\in X_{i}:\|x_{i}\|_{C([0,1])}\leqslant 1\},$
where $\|x\|_{C([0,1])}=\sup_{t\in[0,1]}|x(t)|$ denotes the uniform norm. Hence, each strategy is an entire function rather than a finite-dimensional vector.
The feasible set of player $i$ depends on the opponent’s strategy $x_{-i}$ through
${\bf X}_{i}(x_{-i})=\{\,x_{i}\in X_{i}^{\mathrm{ad}}:x_{i}\neq 0,x_{i}(t)+x_{-i}(t)\leqslant 1\text{ for all }t\in[0,1]\,\}.$
The payoff (or reward) functionals are
$\mathcal{J}_{i}(x_{1},x_{2})=\begin{cases}\displaystyle\int_{0}^{1}tx_{1}(t)dt,&i=1,\\[11.38109pt] \displaystyle\int_{0}^{1}(1-t)x_{2}(t)dt,&i=2.\end{cases}$
Because each player benefits from increasing their function wherever possible, the constraints $x_{i}+x_{-i}\leqslant 1$ and $x_{i}\neq 0$ prevent any pair $(x_{1},x_{2})$ from being mutually optimal. Hence, this game admits no generalized Nash equilibrium. ∎
Notice that $X_{i}^{\rm{ad}}$ in the example above is not compact. Moreover, in infinite dimensions, even compactness can only provide hope that an equilibrium will exist, see [mas-colell_equilibrium_1991] for more examples. In particular, the natural topology in which state equations of PDE-constrained games are well posed is usually only weak or weak-star, hence lower semicontinuity of the induced constraint maps with respect to that topology is often impractical. This motivates the move to structural conditions on the graph, as developed later in Section 5.
@@
We say that the a generalized game $G=({\bf X},\mathcal{J})$ is convex if, for all $i\in I$, the functions ${\bf X}_{i}$ are convex-valued and the functions $\mathcal{J}_{i}(\cdot,x_{-i})$ are convex.
3 Regularity in classical existence theory
This section recalls the standard Kakutani approach to existence of equilibria. The goal is twofold: (i) to pinpoint exactly where lower semicontinuity of the constraint maps enters the argument, and (ii) to set up the notation used later when we replace lower semicontinuity by graph-convexity or the KKM property.
Let $X,Y$ be Banach spaces. The graph of a set-valued map $F:X\to 2^{Y}$ is defined as ${\bf Gr}(F)=\{(x,y)\in X\times Y:y\in F(x)\}$. Further, $F$ is upper semicontinuous at $x_{0}$ if for any open subset $V$ of $Y$ such that $F(x_{0})\subset V$ there exists an open subset $U$ of $X$ containing $x_{0}$ and such that $F(x)\subset V$ for all $x\in U.$ $F$ is lower semicontinuous at $x_{0}$ if for any open subset $V$ of $Y$ such that $F(x_{0})\cap V\neq\emptyset$ there exists an open set $U$ of $X$ containing $x_{0}$ and such that $F(x)\cap V\neq\emptyset$ for all $x\in U.$ Equivalently, $F$ is lower semicontinuous at $x_{0}$ if for any $y_{0}\in F(x_{0})$ and any sequence $x_{n}$ in $X$ converging to $x_{0}$ there exists a sequence $y_{n}$ in $Y$ converging to $y_{0}$ and such that $y_{n}\in F(x_{n}).$ For more details, see [beer_topologies_1993].
Assumption 3.1 (Topology and regularity).
@@ 3 Regularity in classical existence theory
    the map ${\bf X}_{i}:X_{-i}\to 2^{X_{i}}$ is convex-valued and has closed graph.
@@
The assumption above plays a pivotal role in the study of convex GNEPs. Established methodologies, such as the K. Fan inequalities [fan_minimax_1953] and various formulations of Kakutani’s theorem [beer_topologies_1993], rely heavily on this assumption. It is worth mentioning that (ii’) is not actually necessary for any of the results discussed here, but we make it just so there is no need to write $X_{i}^{\rm{ad}}\cap{\bf X}_{i}(x_{-i})$ everywhere.
On the other hand, the assumption below fixes the structure of the objective functions considered in this paper. For further discussion, see [arrow_existence_1954, mas-colell_equilibrium_1991] for a justification of such conditions in the context of economics, for example.
Assumption 3.2 (Objective functions).
For each $i\in I$
@@
    the function $\mathcal{J}_{i}(\cdot,x_{-i}):X_{i}\to\mathbb{R}$ is convex for each $x_{-i}$.
While not the simplest, perhaps the most classical existence result for GNEPs follows from Kakutani’s fixed-point theorem.
Theorem 3.3 (Kakutani’s fixed-point theorem).
Let $K$ be a compact convex subset of a Hausdorff locally convex space and let $F:K\to 2^{K}$ be upper semicontinuous with nonempty, convex, and closed values. Then $F$ has a fixed point.
By sketching the proof of the theorem below, we aim at outlining the standard argument showing how Kakutani’s theorem implies the existence of a Nash equilibrium for a general GNEP and why lower semicontinuity of the constraint maps appears as a crucial assumption.

### v2

@@ Abstract
A generalized Nash equilibrium problem (GNEP) in Banach space consists of $N>1$ optimal control problems with couplings in both the objective functions and, most importantly, in the feasible sets. We address the existence of equilibria for convex GNEPs in Banach space. We show that the standard assumption of lower semicontinuity of the set-valued constraint maps – foundational in the current literature on GNEPs – can be replaced by graph convexity or the so-called Knaster–Kuratowski–Mazurkiewicz (KKM) property. Lower semicontinuity is often essential for obtaining upper semicontinuity of best response maps, crucial for the existence theory based on Kakutani–Fan fixed-point arguments. However, in function spaces or in settings with partial differential equation (PDE) constraints, verifying lower semicontinuity becomes much more challenging (even in convex cases), whereas graph convexity, for example, is often straightforward to check. Our results unify several existence theorems in the literature and clarify the structural role of constraint maps. We also extend Rosen’s uniqueness condition to Banach spaces using a multiplier bias framework.
@@ 1 Introduction
Game theory provides a mathematical framework for studying interactions among decision makers. Games can be cooperative, if enforceable agreements are allowed, or noncooperative [17], and since the seminal work of von Neumann and Morgenstern [36] and Nash’s foundational equilibrium concepts [25, 26], equilibrium analysis has become a central paradigm in economics, engineering, and the social sciences. A generalized Nash equilibrium problem (GNEP) extends the classical setting by allowing each player’s feasible set to depend on the strategies of others. This formulation captures shared or coupled constraints, which arise naturally in models such as electricity markets [21, 22], environmental management [8] and resource allocation [35].
@@
Most classical results on GNEPs are finite-dimensional. However, important applications such as the coupling of energy markets to physical transport of energy carriers (see, e.g., [6] and the references therein) lead to GNEPs with partial differential equations (PDEs) modeling the transport. Such infinite-dimensional formulations, particularly those constrained by PDEs, remain less developed due to compactness and regularity issues. While some works analyze PDE-constrained Nash equilibrium problems via optimality systems [34, 33, 29, 30, 7], genuine generalized frameworks for multiple players have been addressed mainly under jointly convex assumptions [19, 18, 15, 14, 20, 23]. Beyond standard compactness of the feasible strategy sets and closedness of constraint maps, these studies typically require lower semicontinuity of the constraint maps to invoke Kakutani–Fan-type fixed-point theorems. While this is a standard assumption in abstract analysis, it is often not verifiable in function space settings.
@@
This paper contributes to the existence theory for convex GNEPs in Banach spaces by clarifying the trade-off between structural assumptions, such as convexity and regularity of set-valued constraint maps. We show that equilibrium existence can be ensured under graph-convexity or the KnasterKuratowski-Mazurkiewicz (KKM) property without assuming lower semicontinuity. These geometric conditions are often much easier to verify in infinite-dimensional settings, in particular in PDE-constrained contexts and thus broaden the class of problems where equilibrium existence can be rigorously established. This relaxes one of the most restrictive assumptions in the standard theory while retaining the classical framework as a special case.
@@
We also show that Rosen’s uniqueness condition for variational equilibria [32] extends naturally to Banach spaces, and introduce the notion of multiplier bias as a mechanism for equilibrium selection.
@@
Compared with the existing literature on convex GNEPs and quasi-variational inequalities (QVIs) [11, 16, 28, 9], the novelty is not in a single new fixed-point theorem, but in the way how structural assumptions on the constraint maps and the objective functionals can be traded against regularity assumptions, with an eye on PDE-constrained applications. The paper thus complements, rather than overlaps, with the jointly convex theory in [19, 23] and with equilibrium results in function spaces [24].
@@
The rest of the paper is organized as follows: Section 2 introduces the mathematical structure of generalized games in Banach spaces. Section 3 establishes equilibrium existence under graph-convexity and KKM assumptions. Section 4 extends Rosen’s uniqueness result and interprets it as a selection mechanism through multiplier bias.
@@ 2 Generalized finite games in Banach spaces
The strategy set of each player is a fixed Banach space $X_{i}$, likely of infinite dimension, with its topological dual denoted by $X_{i}^{*}$. Representing their private (or individual) constraints, each player observes a nonempty, closed and convex subset $X_{i}^{\rm{ad}}\subset X_{i}$. We write $X\coloneqq X_{1}\times\cdots\times X_{N}$, analogously for $X^{*}$, and $X^{\rm{ad}}\coloneqq X_{1}^{\rm{ad}}\times\cdots\times X_{N}^{\rm{ad}}$ to denote the full strategy space, its topological dual, and the admissible strategy set, respectively.
In game theory, it is common to use the notation $x_{-i}$ to denote the bundle of N1N-1 strategies containing all strategies of a given bundle $x\in X$ except that of player $i.$ Moreover, we write $x=(x_{i},x_{-i})$ when emphasis on player $i$’s strategy is needed, but without changing the original order of the components of $x$. Consequently, we let $X_{-i}\coloneqq X_{1}\times\cdots\times X_{i-1}\times X_{i+1}\times\cdots\times X_{N}$ (and the corresponding $X_{-i}^{\rm{ad}}$) and say that $x_{-i}\in X_{-i}.$ In earlier literature on Nash games the bundle $x_{-i}$ is sometimes referred to as an $i$-incomplete combination [17].
@@
$\begin{cases}\text{Given }x_{-i}\in X_{-i}^{\rm{ad}};\\ \text{Minimize }\mathcal{J}_{i}(x_{i},x_{-i})\ \text{over }x_{i}\in X_{i}^{\rm{ad}}\cap{\bf X}_{i}(x_{-i}).\end{cases}$ ($P_{i}$)
@@
As a first example, the simplest situation in which generalized games appear is when a constraint is shared among players. This is the case when, for example, a regulatory individual or agency determines that the strategy bundle $x$ must belong to a certain closed set $\mathcal{C}\subset X.$ This leads to non-constant constraint maps
@@
where $\pi_{i}$ denotes the projection onto $X_{i}$.
We are interested in the concept of Nash equilibrium [25, 26] as the notion of solution for a generalized game.
@@
Very simple finite dimensional examples can illustrate the fact that constant versus non-constant constraint maps may drastically change the set of equilibria, see e.g. [11]. Also, the issue of compacteness (of the feasible strategy sets) becomes essential in games with nonsmooth objectives [24]. Furthermore, in infinite dimensions, even compactness cannot ensure that an equilibrium exists, see [24] for examples.
In particular for PDE-constrained games, the natural topology in which state equations are well posed, closed and bounded sets are usually only weak or weak-star compact, hence lower semicontinuity of the induced constraint maps with respect to that topology is often impractical. This motivates the move to structural conditions on the graph, as we propose in this paper.
@@
We say that the generalized game $G=({\bf X},\mathcal{J})$ is convex if, for all $i\in I$, the functions ${\bf X}_{i}$ are convex-valued and the functions $\mathcal{J}_{i}(\cdot,x_{-i})$ are convex.
Let $X,Y$ be Banach spaces. The graph of a set-valued map $F:X\to 2^{Y}$ is defined as ${\bf Gr}(F)=\{(x,y)\in X\times Y:y\in F(x)\}$. We say that $F$ is lower semicontinuous (l.s.c) at $x_{0}$ if for any open subset $V$ of $Y$ such that $F(x_{0})\cap V\neq\emptyset$ there exists an open set $U$ of $X$ containing $x_{0}$ and such that $F(x)\cap V\neq\emptyset$ for all $x\in U.$ Equivalently, $F$ is l.s.c. at $x_{0}$ if for any $y_{0}\in F(x_{0})$ and any sequence $x_{n}$ in $X$ converging to $x_{0}$ there exists a sequence $y_{n}$ in $Y$ converging to $y_{0}$ and such that $y_{n}\in F(x_{n}).$ $F$ is said to be l.s.c on a subset $\mathcal{X}$ of $X$ if it is l.s.c at all $x_{0}\in\mathcal{X}.$ For more details, see [4].
Along with lower semicontinuity of the constraint maps ${\bf X}_{i}$ (for each $i$), the following two assumptions are standard in the study of GNEPs.
Assumption 2.4.
@@
    the map ${\bf X}_{i}:X_{-i}\to 2^{X_{i}}$ is convex-valued and has closed graph;
@@
Assumption 2.5.
For each $i\in I$ and $x_{-i}\in X_{-i}$
@@
    the function $\mathcal{J}_{i}(\cdot,x_{-i}):X_{i}\to\mathbb{R}$ is convex.
Assumption 2.4 above plays a pivotal role in the study of convex GNEPs. Established methodologies, such as the K. Fan inequalities [13] and various formulations of Kakutani’s theorem [4], rely heavily on this assumption. It is worth mentioning that (ii’) is not actually necessary for any of the results discussed here, but we make it just so there is no need to write $X_{i}^{\rm{ad}}\cap{\bf X}_{i}(x_{-i})$ everywhere.
On the other hand, Assumption 2.5 fixes the structure of the objective functions considered in this paper. See, for example, [3, 24] for a justification of these conditions in the context of economics.
Assuming lower semicontinuity of ${\bf X}_{i}$ is crucial for the closedness of the so called graph of the best-response map. Fixed points of such maps characterize Nash equilibria and this is why Kakutani’s theorem is central in the current existence theory for GNEPs. This assumption is, however, rarely easy to verify in practical settings such as PDE-constrained or function space games, see [5, 2, 11, 12]. This limitation partly explains the popularity of models where the coupling constraints have the structured form (2.2), for which the lower semicontinuity issue can be bypassed by exploiting properties of the Nikaido-Isoda function. This shows that to a certain extent, regularity (lower semicontinuity) can be traded by structure (jointly convex games).
Our main result in this paper extends this idea to a general class of convex GNEPs (not only jointly convex). It says that if, in addition to convexity of GNEP, the graph of the bundle constraint map $\bf X$}, i.e. the set $\{(x,y)\in X\times X:y\in{\bf X}(x)\}$ is convex, then lower semicontinuity of the maps ${\bf X}_{i}$ is no longer necessary for existence of an equilibrium. Details on this are discussed in the next section.
3 Potential lack of lower semicontinuity of constraint maps
For convex games we show next that lower semicontinuity of the constraint maps is not needed for equilibrium existence once additional geometric structure is available. This allows us to cover games with significantly more complex and heterogeneous constraints, such as generalized PDE-constrained games where each player’s state equation depends on forecasts of incomplete or partially available information.
Definition 3.1.